OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (1, 5) and (-2, 3)

OpenStudy (amistre64):

slope is defined as the ration of how far y moves with respect to x y moves from 5 to 3 which is a distance of -2 as: x moves from 1 to -2 which is a distance of -3 slope = change in y/change in x slope = 2/3. ax+by = n is the standard form of the equation where slope = -a/b -2x +3y = n ; now calibrate it with either point; i pick (1,5) -2(1) +3(5) = n -2 +15 = n = 13 -2x +3y = -15 is good, but some dont like a negative in front ... 2x -3y = 15

OpenStudy (amistre64):

try that one more time .... -2x +3y = 13 2x -3y = -13 and thats my final offer :)